A 60kg hiker climbs up to the top of a hill that is 250m vertically above the starting position. The change is gravitational potential energy is
1
Expert's answer
2020-04-29T09:49:03-0400

potential energy down the mountain

"M_0= m \cdot g \cdot h_0=60 \cdot 9.81 \cdot 0 =0"

potential energy on top of a mountain

"M= m \cdot g \cdot h=60 \cdot 9.81 \cdot 250 =1.471 \cdot 10^5 J"

Change in gravitational potential energy

"\Delta M= M-M_0=1.471 \cdot 10^5-0=1.471 \cdot 10^5J"